OpenStudy (anonymous):

How far up a wall will an 11 meter ladder reach, if the foot of the ladder must be 4 meters from the base of the wall? Round to the nearest hundredth.<-- help please

OpenStudy (anonymous):

use pyrhagoras theorem 11^2 = 4^2 + h^2 h^2 = 121- 16 = 105 h = sqrt 105 = 10.25 meters

OpenStudy (anonymous):

thank you